OpenStudy (anonymous):

medal and fan Which equation would best help solve this problem? Five added to 3 times a number is equal to 8 less than five times the number. (Points : 1) 3x + 5 = 5x – 8 3x – 4 = 5x + 8 4x + 3 = 8x – 5 8x + 5 = 4x + 3

OpenStudy (anonymous):

Five added to 3 times a number is equal to 8 less than five times the number 5 + (3*x) = -8 + (5*x) move the terms 3x +5 = 5x -8